OpenStudy (anonymous):

sin(20-π) = sinΘ SOLVE ON THE INTERVAL: 0≤Θ<2π

OpenStudy (anonymous):

@mathstudent55 help pls?

OpenStudy (anonymous):

use the law \[\sin \left( x-y \right)=\sin x \cos y-\cos x \sin y\]

OpenStudy (anonymous):

ok then..?

OpenStudy (anonymous):

then u have the value of \[\sin \left( 20-\Pi \right)\]

OpenStudy (anonymous):

achieve it first

OpenStudy (anonymous):

but whgats sin20

OpenStudy (anonymous):

are u have a calculator?

OpenStudy (anonymous):

yea but i mean find exact value? shouldnt in be in like radical form?

OpenStudy (anonymous):

for example: sqrt 3 /2 ?

OpenStudy (anonymous):

wait never mind.. i get that part

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait but how do you get sin 20? cuz you cant use the sin(A-B) rule cuz theres nothing that willsubtract to get 20....

OpenStudy (anonymous):

@amistre64